OpenStudy (anonymous):

Simplify. Assume that no denominator is equal to zero. ([3^2]^3g^5h^8)^2

OpenStudy (anonymous):

\[([3^{2}]^{3}g ^{5}h ^{8})^{2}\]

OpenStudy (anonymous):

Lets start with the\[(3^2)^3\]\[(x^m)^n=x^{mn}\]So what do you think this is?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

so it would be 729g^10h^16?

hartnn (hartnn):

\(g^{10}h^{16}\) part is correct. what about ? \(([3^2]^3)^2\)=? its not 729

hartnn (hartnn):

and it would be better if u keep it in power of 3, \(3^{?}\)

OpenStudy (anonymous):

531,441?

hartnn (hartnn):

yes. if u need to keep it in exponent form, it will be \(\huge 3^?\)

OpenStudy (anonymous):

okay thanks!

hartnn (hartnn):

\(([3^2]^3)^2=3^{12}\) welcome.